Juice comes in cylindrical cans of two sizes. The larger can has a radius which is three times the radius of the smaller can. Both cans have the same height. How much more juice is in the large can than the small can? f. 9 times as much g. 6 times as much h. 3 times as much j. 27 times as much


Answer by mananth(16949)

Juice comes in cylindrical cans of two sizes. The larger can has a radius which is three times the radius of the smaller can. Both cans have the same height. How much more juice is in the large can than the small can? f. 9 times as much g. 6 times as much h. 3 times as much j. 27 times as much

Let the radius of the smaller can be x
the larger can radius = 3x
volume of smaller can = pi*x^2 *h
volume of bigger can = pi* (3x)^2 *h
Thpi*(3x)^2*h / pi*x^2*h
the ratio of both
= pi*(3x)^2*h / pi*x^2*h
=pi*9x^2*h / pi*x^2*h
= 9
So the larger can contains 9 times the quantity of the smaller can